# A Complete Active Space Self-Consistent Field (CASSCF) approach for molecules in QED environments

### v1

@@ Abstract
Multireference systems are usually challenging to investigate using ab initio methods as they require an accurate description of static electron correlation. The urgency of developing similar approaches is even more pressing when molecules strongly interact with light in quantum-electrodynamics (QED) environments. In fact, in this context, multireference effects might be induced or reduced by the presence of the field. In this work, we extend the Complete Active Space Self-Consistent Field (CASSCF) approach to polaritonic systems. The method is tested on benchmark multireference problems and applied to investigate field-induced effects on the electronic structure of well-known multiconfigurational processes. We analyze the strengths and limitations of the method with particular attention to the appearance of possible origin dependencies.
Perugia] Dipartimento di Chimica, Biologia e Biotecnologie, Università degli Studi di Perugia, Via Elce di Sotto, 8,06123, Perugia, Italy Trondheim] Department of Chemistry, Norwegian University of Science and Technology, 7491 Trondheim, Norway Perugia] Dipartimento di Chimica, Biologia e Biotecnologie, Università degli Studi di Perugia, Via Elce di Sotto, 8,06123, Perugia, Italy
@@ 1 Introduction
Manipulation of molecular properties by quantum fields is becoming nowadays a very promising non-intrusive alternative to chemical controls 1, 2, 3, 4, 5, 6, 7, 8, 9. The observation of these effects was possible thanks to the significant technological advances made in the fabrication of efficient optical devices (optical 10, 11, plasmonic 12, 13 and superconducting 14, 15, 16 cavities, waveguides 17, 18, 19, etc.). The simplest of these devices is the Fabry-Pèrot optical cavity, device composed of two mirrors whose separation, L, is related to the wavelength of the field inside the cavity. In these conditions, energy is exchanged between the system and the cavity field, inducing a phenomenon called quantum Rabi oscillations. Due to the strong light-matter interaction, photons and molecular states lose their individuality and show mixed molecular and photonic character, generating hybrid states known as ”polaritons” 20, 21, 22, 23. The formation of polaritons is associated with a splitting of the energy levels, known as Rabi splitting (), that increases with the strength of the light-matter coupling () (see Fig. 1).
@@
It is noteworthy to point out that molecules can interact strongly with the field also in the absence of real photons, through a direct coupling between the molecular states and the vacuum fluctuations of the electromagnetic field. In the past decade, a lot of research was devoted to explore the potential of strong light-matter interaction on a plethora of physical or chemical phenomena. In particular, by coupling molecules or materials with different optical devices it was possible to manipulate the absorption, emission, and scattering properties of molecules and materials 24, 25, 26, 27, to induce Bose–Einstein condensation of polaritons 28, 29, 30, to increase conductivity in organic semiconductors 31, or to manipulate the spin properties of matter 32, 33, 34. One of the most debated but also interesting aspects was observed in chemistry, where field-induced effects on chemical reactivity were demonstrated to inhibit, catalyze, and even make some reactions selective toward specific products 35, 36, 37, 38.
@@
Despite the large attention devoted to this field in the past decade, the physics behind many of these processes still remains unclear, and progresses are still very slow due to the many limitations in the validation of the experimental setup 39, 40, 41. For these reasons, the development of effective theoretical approaches is crucial to overcome experimental limitations and to propose a consistent description of these phenomena 42, 43. In recent years, a small number of research groups started developing some extensions to the most commonly used ab initio methodologies, i.e., Density Functional Theory (DFT) 44, 45, 46, Hartree-Fock (HF), Full Configuration Interaction (FCI), Coupled Cluster (CC) 47, 48, 49, 50, 51, and second-order Møller-Plesset perturbation theory (MP2) 52, to investigate coupled electron-photon systems. Recently, fully relativistic versions of QED-HF and QEDFT have been also proposed to investigate molecular systems containing heavy atoms in strong coupling conditions 53, 54. The quantum electrodynamics (QED) extensions to HF and DFT are quite efficient and can be easily applied to relatively large systems. However, they completely miss correlation or describe it in a very approximated way. Moreover, in the DFT case, an effective exchange-correlation functional for the electron-photon interaction is still far from being formulated 42.
@@
Coupled Cluster theory is known to be the most reliable method in quantum chemistry to simulate molecular systems and also for interacting electron-photon systems, its QED extension is at the moment the reference ab-initio methodology. However, as for the standard electronic theory, the applicability of this method is limited to small to medium sized molecules and its description becomes qualitatively poor when multiconfigurational effects come into play. This problem is well-known and it prevents the investigation of a relatively large class of complexes, including several systems involved in chemically and biologically interesting precesses 55. Belong to these systems, open-shell transition metal complexes, (poly)radicalic systems, excited states of organic molecules, and torsion or breakings of conjugated bonds 56, 57.
@@
The state-of-the-art reference approaches usually used to investigate these systems are the Multiconfigurational Self-Consistent Field (MCSCF) methods, which are based on a wavefunction defined as linear combination of a reduced number of determinants or Configuration State Functions (CSFs). Nowadays a wide variety of MCSCF methods have been developed. The main difference between them lies in the protocol employed to select the determinants to include in the expansion 58. The most popular MCSCF method is the Complete Active Space Self-Consistent Field (CASSCF), which is based on a partitioning of the orbital space into three classes: inactive, virtual, and active orbitals. The multideterminantal expansion in the case of CASSCF theory is obtained by choosing a certain number of so-called active electrons and computing all the possible excitations of these electrons within the active space. If the active space is chosen properly, (i.e., the most correlated orbitals and electrons are included)the CASSCF wavefunction is able to provide a qualitatively-correct description of the molecular system. Despite this method is equivalent to a FCI expansion within the active space, with the current computational facilities, it can be applied, in a relatively efficient way, to active spaces with up to 20 electrons in 20 orbitals 59.
@@
In the latest years, some of the most common ab-initio multireference methods have been extended to coupled-light matter systems 60, 61, 62. In particular, the implementation of QED-CASCI and QED-DMRG provided the first steps towards the inclusion of static correlation effects in optical cavities. However, the CASCI procedure does not optimize the orbitals, leading to variationally higher energies with respect to its SCF counterpart 63, and although the DMRG method has a polynomial scaling, its cost is higher than CASSCF for small active spaces.
In this work, we present the first extension of CASSCF theory to coupled light-matter systems (QED-CASSCF). The implementation is performed following a restricted-step second order optimization scheme, according to the implementations in Refs. 64, 65, 66, 67.
@@
This paper is organized as follows: in Section 2, a brief summary of the CASSCF method and its second-order implementation is presented. In Section 3, the fundamental details of QED-CASSCF are described. In Section 4, QED-CASSCF is tested on small molecules against other ab-initio methodologies. In this section, QED-CASSCF is also used to investigate the field effects on well known multireference chemical processes. We conclude the work with future perspectives presented in Section 5.
@@ 2 CASSCF Theory
and it is optimized by variationally minimizing the energy with respect to the molecular orbital (MO) and configuration interaction (CI) coefficients ( and , respectively)
Among the MCSCF methods 68, 69, 70, 71, 72, the Complete Active Space Self-Consistent Field (CASSCF) approach is still the state-of-the art procedure for molecular systems dominated by static correlation. The orbital space of a molecule in the CASSCF approximation is partitioned into three subspaces, as depicted in Fig. 2: inactive orbitals are always fully occupied, active orbitals have non-integer occupations, and virtual orbitals are always empty. In this work the different orbitals are addressed according to the following indices choice: for inactive orbitals, for active orbitals, for virtual orbitals, and for generic orbitals.
The multideterminantal expansion for the wavefunction is obtained considering all the determinants computed by distributing active electrons in active orbitals, namely CAS(N,M).
The first implementations of CASSCF relied on first-order solvers and among these, the super-CI approach is the most widely applied 73, 74. However, using this scheme, the optimization of the CASSCF wavefunction is affected by severe convergence issues. This led to the development of robust second-order quadratically-convergent solvers which exhibit better convergence properties in a reduced number of iterations 64, 75, 76, 77, 78, 79, 80, 81, 82, 83, 67. CASSCF implementations based on second-order algorithms are nowadays relatively well-established in the quantum chemistry community and are available in the majority of the mostly used software packages 84, 85, 86, 87, 88. This approach is particularly convenient for the application to simulate small molecular systems, and calculations on larger systems are nowadays efficiently performed with the aid of parallelization, AO-based implementations, or by exploiting the Cholesky decomposition or the resolution of identity (RI) approximation of the two-electron integrals 83, 67.
@@
where is the spin traced singlet excitation operator. While no alternatives to the parametrization of the orbital space are available, the CI space is liable to multiple choices depending on the CASSCF variant that has to be implemented. In the case of State-Specific (SS) CASSCF method, a linear parametrization for the CI part is commonly adopted and the wavefunction can be written as 64:
where is indicated as the current expansion point (CEP) and constitutes the current approximation to the wavefunction. , instead, is a projector operator that removes the components parallel to the CEP from .
The choice of the exponential parametrization allows to consider only the non-redundant rotations between the orbital spaces (i.e., orbital rotations which contribute to the energy: inactive-active, inactive-virtual, and active-virtual). The linear parametrization of the CI space contains only one redundant parameter when is parallel to the reference state . However, this is not an issue as this redundancy can be easily tracked and it does not interfere with the optimization procedure.
In the following of this section we will quickly recap the implementation details of the second-order SS-CASSCF algorithm that will be applied, in the next section, for the extension to strongly coupled electron-photon systems.
@@ 2.1 Second-Order CASSCF
where indicates a small variation of the wavefunction parameters (, ). In Eq. 5, is the CASSCF energy at the current expansion point. is the gradient vector containing the first derivatives of the energy with respect to the configuration (c) and orbital (o) coefficients calculated at the current expansion point:
@@
Also the Hessian matrix elements are calculated at the current expansion point.
@@
Here, are the so-called trial vectors used in the Davidson scheme 89 which contain the wavefunction parameters with respect to the orbital and configuration parts
This approach allows for large scale operations without explicitly calculating and storing the blocks of in memory 90. Moreover, sigma vectors can be conveniently computed applying minor modifications to intermediate quantities already calculated for the gradients.
@@
This procedure is not optimal since it requires the Hessian to be positive-definite. This condition is not always guaranteed, especially at the beginning of the optimization procedure. Even when it is respected, the computed step may not point in the direction of the minimum, leading to a poor convergence 91. This problem can be largely reduced by adopting a restricted step procedure 64, 92. According to this scheme, the choice of the step is restricted within a well-defined trust region with radius, :
where we assume the potential energy surface to be correctly approximated by the quadratic model. This ensures the Hessian to be positive-definite. This procedure was first described by Levenberg and Marquardt 93, 94 and is based on the solution of the level-shifted Newton equation:
@@
The relation between the eigenvalues of and is given by the Hylleraas-Undheim-MacDonald theorem 95, 96:
@@
where is the eigenvalue of the Hessian and is a scaling parameter that forces the step to lie withing the trust region. The trust radius is changed adaptively during the procedure following the strategy described by Fletcher in Ref. 91. As long as the step is kept within the trust region, the Newton’s step in 10 is recovered and convergence to the closest minimum is guaranteed.
@@
Calculation of the gradient by Eq. 6. If its norm is lower than a certain threshold, convergence is reached;
@@
Check the step and adjust the trust radius according to the procedure described by Fletcher in Ref. 91;
@@ 3 QED-CASSCF
The description of molecular systems in optical cavities requires to account explicitly for the quantum character of the electromagnetic field. In our work we will use the well-established non-relativistic Pauli-Fierz Hamiltonian in length gauge and dipole approximation as starting point for the development of the QED-CASSCF approach 47:
Here is the dipole moment operator, is its expectation value and is the light-matter coupling parameter that depends on the quantization volume and the polarization vector . For simplicity a single cavity mode is included in our treatment, but extension to more modes can be applied in a trivial manner. To ensure origin invariance, Hamiltonian 16 has been already expressed in the coherent-state basis 47. Following the notation used in Ref. 98, we can conveniently set and rewrite the Pauli-Fierz Hamiltonian as
In this form the dipole-self energy term is included into the one- and two-electron integrals
@@
For a general QED-MCSCF theory the wavefunction can be expressed as 60:
@@
In this way a generic polaritonic state can be written as
with representing the ground state.
For the sake of readability of the following section we conveniently separate the Pauli-Fierz Hamiltonian into three terms:
@@
The extension of CASSCF to molecular polaritons and its second-order implementation requires the computation of modified energy contributions, gradients and sigma vectors. The expression for the QED-CASSCF energy can be written as:
@@
has the same form of the purely electronic CASSCF energy 83 with
being the one- and two-body reduced density matrices, respectively, and
the inactive Fock matrix. The purely photonic contribution to the energy is given by:
@@
is made of two ”off-diagonal” contributions
and
The general expression for the orbital part of the gradient is given by:
The purely electronic contribution is known 58:
@@
is the generalized Fock matrix. It is easily verified that the purely photonic contribution to the orbital gradient is zero since and commute. Finally, the interaction term to the gradient reads as:
where the one-body density matrices are defined as:
Finally, the sigma vectors defined in Eq. 8 can be calculated as:
where is the one-index transformed Hamiltonian defined as 58:

### v3

# A Complete Active Space Self-Consistent Field approach for molecules in QED environments

@@ Abstract
Multireference systems are usually challenging to investigate using ab-initio methods as they require an accurate description of static electron correlation. The urgency of developing similar approaches is even more pressing when molecules strongly interact with light in quantum-electrodynamics (QED) environments. In fact, in this context, multireference effects might be induced or reduced by the presence of the field. In this work, we extend the Complete Active Space Self-Consistent Field (CASSCF) approach to polaritonic systems. The method is tested on benchmark multireference problems and applied to investigate field-induced effects on the electronic structure of well-known multiconfigurational processes. Strengths and limitations of the method have been thoroughly analyzed.
Perugia] Dipartimento di Chimica, Biologia e Biotecnologie, Università degli Studi di Perugia, Via Elce di Sotto, 8,06123, Perugia, Italy Trondheim] Department of Chemistry, Norwegian University of Science and Technology, 7491 Trondheim, Norway Trondheim] Department of Chemistry, Norwegian University of Science and Technology, 7491 Trondheim, Norway Perugia] Dipartimento di Chimica, Biologia e Biotecnologie, Università degli Studi di Perugia, Via Elce di Sotto, 8,06123, Perugia, Italy
@@ 1 Introduction
Manipulation of molecular properties by quantum fields is becoming nowadays a very promising non-intrusive alternative to chemical control 1, 2, 3, 4, 5, 6, 7, 8. The observation of these effects was possible thanks to the significant technological advances made in the fabrication of efficient optical devices (optical 9, 10, plasmonic 11, 12 and superconducting 13, 14, 15 cavities, waveguides 16, 17, 18, etc.). The simplest of these devices is the Fabry–Pérot optical cavity (see Fig. 1), a device composed of two mirrors whose separation, L, is related to the fundamental wavelength of the field inside the cavity. In these conditions, energy is coherently exchanged between the molecule and the cavity field, inducing a phenomenon called Rabi oscillations. Due to the strong light-matter interaction, photons and molecular states lose their individuality and show mixed molecular and photonic character, generating hybrid states known as ”polaritons” 19, 20, 21, 22. The formation of polaritons is associated with a splitting of the energy levels, known as Rabi splitting (), that increases with the strength of the light-matter coupling ().
@@
It is noteworthy to point out that molecules can interact strongly with the field also in the absence of real photons, through a direct coupling between the molecular states and the vacuum fluctuations of the electromagnetic field. In the past decade, a lot of research was devoted to exploring the potential of strong light-matter interaction on a plethora of physical or chemical phenomena. In particular, by coupling molecules or materials with different optical devices, it was possible to manipulate the absorption, emission, and scattering properties of molecules and materials 23, 24, 25, 26, to induce Bose–Einstein condensation of polaritons 27, 28, 29, to increase conductivity in organic semiconductors 30, or to manipulate the spin properties of matter 31, 32, 33. One of the most debated but also interesting aspects was observed in chemistry, where field-induced effects on chemical reactivity were demonstrated to inhibit, catalyze, and even make some reactions selective toward specific products 34, 35, 36, 37.
@@
Despite the large attention devoted to this field in the past decade, the physics behind many of these processes still remain unclear, and progress is still slow due to the many limitations in the validation of the experimental setup 38, 39, 40. For these reasons, the development of effective theoretical approaches is crucial to overcome experimental limitations and to propose a consistent description of these phenomena 41, 42. In recent years, a small number of research groups started developing some extensions to the most commonly used ab-initio methodologies, i.e., Density Functional Theory (DFT) 43, 44, 45, Hartree-Fock (HF), Full Configuration Interaction (FCI), Coupled Cluster (CC) 46, 47, 48, 49, 50, and second-order Møller-Plesset perturbation theory (MP2) 51, 52, to investigate coupled electron-photon systems. Recently, fully relativistic versions of QED-HF and QEDFT have also been proposed to investigate molecular systems containing heavy atoms in strong coupling conditions 53, 54. The quantum electrodynamics (QED) extensions to HF and DFT are quite efficient and can be easily applied to relatively large systems. However, they completely miss correlation or describe it in a very approximate way. Moreover, in the DFT case, an effective exchange-correlation functional for the electron-photon interaction is still far from being formulated 41.
@@
Coupled Cluster theory is known to be the most reliable method in quantum chemistry to simulate molecular systems and also for interacting electron-photon systems, its QED extension is at the moment the reference ab-initio methodology. However, as for the standard electronic theory, the applicability of this method is limited to small- to medium-sized molecules, and its description becomes qualitatively poor when multiconfigurational effects come into play. This problem is well-known and it prevents the investigation of a relatively large class of complexes, including several systems involved in chemically and biologically interesting processes 55. For example, relevant molecular systems that can hardly be described with CC methodologies include open-shell transition metal complexes, (poly)radicalic systems, excited states of organic molecules, and torsion or breakings of conjugated bonds 56, 57.
@@
The state-of-the-art reference approaches usually used to investigate these systems are the Multiconfigurational Self-Consistent Field (MCSCF) methods, which are based on a wavefunction defined as a linear combination of a reduced number of determinants or Configuration State Functions (CSFs). A wide variety of MCSCF methods have been developed. The main difference between them lies in the protocol employed to select the determinants to include in the expansion 58. The most popular MCSCF method is the Complete Active Space Self-Consistent Field (CASSCF), which is based on a partitioning of the orbital space into three classes: inactive, virtual, and active orbitals. The multideterminantal expansion in the case of CASSCF theory is obtained by choosing a certain number of so-called active electrons and computing all the possible excitations of these electrons within the active space. If the active space is chosen properly (i.e., the most correlated orbitals and electrons are included), the CASSCF wavefunction is able to provide a qualitatively-correct description of the multireference molecular system. Despite this method being equivalent to a FCI expansion within the active space, with the current computational facilities, it can be applied, in a relatively efficient way, to active spaces with up to 20 electrons in 20 orbitals 59.
@@
In the latest years, some of the most common ab-initio multireference methods have been extended to coupled light-matter systems 60, 61, 62. In particular, the implementation of QED-CASCI and QED-DMRG provided the first steps towards the inclusion of static correlation effects in optical cavities. However, the CASCI procedure does not optimize the orbitals, leading to variationally higher energies with respect to its SCF counterpart 63, and although the DMRG method has a polynomial scaling, its cost is higher than CASSCF for small active spaces. The comparatively higher computational cost of DMRG for small active spaces provides a rationale for extending and applying the CASSCF method to the QED case.
In this work, we present an extension of the State-Specific (SS-)CASSCF theory to coupled light-matter systems (QED-CASSCF). The implementation is performed following a restricted-step second-order optimization scheme, according to the implementations in Refs. 64, 65, 66, 67. Recently, a QED version of the State-Averaged (SA) variant of CASSCF was proposed by Vu et al. 68 in parallel to this work. It was applied to a small test case with an active space close to the full orbital space, with results very close to the FCI limit. In this work, we perform simulations of larger systems where the CASSCF method does not approach the FCI limit. In this situation, a deep analysis of the method’s performances to describe field-induced effects on multireference systems could be performed.
@@
This paper is organized as follows: in Section 2, a brief summary of the CASSCF method and its second-order implementation is presented. In Section 3, the fundamental details of QED-CASSCF are described. In Section 4, QED-CASSCF is tested on small molecules against other ab-initio methodologies. In this section, QED-CASSCF is also used to investigate the field effects on well-known multireference chemical processes. We conclude the work with future perspectives presented in Section 5.
@@ 2 CASSCF Theory
and it is optimized by variationally minimizing the energy with respect to the molecular orbital-rotation parameters and configuration interaction (CI) coefficients ( and , respectively)
Among the MCSCF methods 69, 70, 71, 72, 73, the Complete Active Space Self-Consistent Field (CASSCF) approach is still the state-of-the-art procedure for molecular systems dominated by static correlation. The orbital space of a molecule in the CASSCF approximation is partitioned into three subspaces, as depicted in Fig. 2: inactive orbitals are always fully occupied, active orbitals have non-integer occupations, and virtual orbitals are always empty. In this work the different orbitals are addressed according to the following indices choice: for inactive orbitals, for active orbitals, for virtual orbitals, and for generic orbitals.
The multideterminantal expansion for the wavefunction is obtained by considering all the determinants computed by distributing active electrons in active orbitals, namely CAS(N,M).
The first implementations of CASSCF relied on first-order solvers, and among these, the super-CI approach is the most widely applied 74, 75. Although recent first-order implementations proved to be very efficient 76, during the optimization of the CASSCF wavefunction, convergence issues may arise. This led to the development of robust second-order quadratically-convergent solvers, which exhibit better convergence properties in a reduced number of iterations at the price of a higher cost for each step 64, 77, 78, 79, 80, 81, 82, 83, 84, 85, 67. CASSCF implementations based on second-order algorithms are nowadays relatively well-established in the quantum chemistry community and are available in the majority of the most used software packages 86, 87, 88, 89, 90. This approach is particularly convenient for the application to simulate small molecular systems, and calculations on larger systems are nowadays efficiently performed with the aid of parallelization, AO-based implementations, or by exploiting the Cholesky decomposition or the resolution of identity (RI) approximation of the two-electron integrals 85, 67.
@@
where is the spin traced singlet excitation operator with . While no alternatives to the parametrization of the orbital space are available, the CI space is liable to multiple choices depending on the CASSCF variant that has to be implemented. In the case of State-Specific (SS) CASSCF method, a linear parametrization for the CI part is commonly adopted and the wavefunction can be written as 64:
where is indicated as the current expansion point (CEP) and constitutes the current approximation to the wavefunction. , instead, is a projector operator that removes the components parallel to the CEP from the correction vector, , which contains the variational parameters, , with respect to the CI part:
The choice of the exponential parametrization allows to consider only the non-redundant rotations between the orbital spaces (i.e., orbital rotations that contribute to the energy: inactive-active, inactive-virtual, and active-virtual) 58. The linear parametrization of the CI space contains only one redundant parameter, that is, when the correction vector has components parallel to the reference state . However, this is not an issue, as this redundancy can be easily tracked and is projected out by the operator . Therefore, this redundancy does not interfere with the optimization procedure.
In the following of this section, we will quickly recap the implementation details of the second-order SS-CASSCF algorithm that will be applied, in the next section, for the extension to strongly coupled electron-photon systems.
@@ 2.1 Second-Order CASSCF
where indicates a small variation of the wavefunction parameters (, ). In Eq. 6, is the CASSCF energy and is the gradient vector containing the first derivatives of the energy with respect to the configuration (c) and orbital (o) coefficients, all calculated at the current expansion point:
@@
The Hessian matrix elements are calculated, again, at the current expansion point.
@@
Here, are the so-called trial vectors used in the Davidson scheme 91 which contain the wavefunction parameters with respect to the orbital and configuration parts
This approach allows for large scale operations without explicitly calculating and storing the blocks of in memory 92. Moreover, sigma vectors can be conveniently computed applying minor modifications to intermediate quantities already calculated for the gradients.
@@
This procedure is not optimal since it requires the Hessian to be positive-definite. This condition is not always guaranteed, especially at the beginning of the optimization procedure. Even when it is respected, the computed step may not point in the direction of the minimum, leading to a poor convergence 93. This problem can be largely reduced by adopting a restricted step procedure 64, 94. According to this scheme, the choice of the step is restricted within a well-defined trust region with radius, :
where we assume the potential energy surface to be correctly approximated by the quadratic model. This ensures the Hessian to be positive-definite. This procedure was first described by Levenberg and Marquardt 95, 96 and is based on the solution of the level-shifted Newton equation:
@@
The relation between the eigenvalues of and of is given by the Hylleraas-Undheim-MacDonald theorem 97, 98:
@@
where is the eigenvector of the Hessian and is a scaling parameter that forces the step to lie withing the trust region. The trust radius is changed adaptively during the procedure following the strategy described by Fletcher in Ref. 93. As long as the step is kept within the trust region, the Newton’s step in 11 is recovered, and convergence to the closest minimum is guaranteed.
@@
Compute the gradient in Eq. 7. If its norm is lower than a certain threshold, convergence is reached;
@@
Check the step and adjust the trust radius according to the procedure described by Fletcher in Ref. 93;
@@ 3 QED-CASSCF
The description of molecular systems in optical cavities requires to account explicitly for the quantum character of the electromagnetic field. In our work, we will use the well-established non-relativistic single-mode Pauli-Fierz Hamiltonian in length gauge and dipole approximation as a starting point for the development of the QED-CASSCF approach 46:
Here, is the dipole moment operator, is its expectation value, and is the light-matter coupling parameter that depends on the quantization volume and the polarization vector . For simplicity, a single cavity mode is included in our treatment, but extension to more modes can be applied in a trivial manner. To ensure origin invariance, the Hamiltonian in Eq. 17 has already been expressed in the coherent-state basis 46. Following the notation used in Ref. 100, we can conveniently set and rewrite the Pauli-Fierz Hamiltonian as
In this form the dipole-self energy term is included in the one- and two-electron integrals
@@
For a general QED-MCSCF theory, the ground-state wavefunction can be expressed as 60:
@@
For the sake of readability of the following section, we conveniently separate the Pauli-Fierz Hamiltonian into three terms:
@@
The extension of CASSCF to molecular polaritons and its second-order implementation requires the computation of modified energy contributions, gradients, and sigma vectors. The expression for the QED-CASSCF energy can be written as:
@@
Are the one- and two-body reduced density matrices, respectively, where
is the two-electron excitation operator running on the active indices 58. In Eq. 24
are the elements of the inactive Fock matrix 77, 58, and
is denoted as inactive energy 85. The purely photonic contribution to the energy is given by:
@@
where the modified one-body density matrix is defined as:
@@
The general expression for the orbital part of the gradient from Eq. 7 is given by:
where the purely photonic contribution to the orbital gradient is zero since and commute. The purely electronic contribution is known 58:
@@
is the generalized Fock matrix. Finally, the interaction term to the gradient reads as:
The sigma vectors defined in Eq. 9 can be calculated as:
where we have introduced the modified state
is the one-index transformed Hamiltonian 58, where only the terms involving molecular integrals are one-index transformed:
@@
Substituting these quantities in Eqs. 7-9 an efficient second-order algorithm for QED-CASSCF can be naturally obtained.